OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work. Write your final answer exactly as a fraction in lowest terms. 1/11+5/6=

sam (.sam.):

\[\frac{1}{11}+\frac{5}{6}\] First make denominator the same \[\frac{1 \times 6}{11\times 6}+\frac{5\times 11}{6 \times 11}\] \[\frac{6}{66}+\frac{55}{66}\] Once the denominator its same, you can add the fraction together \[\frac{6+55}{66}\] \[\frac{61}{66}\]